OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4.

jimthompson5910 (jim_thompson5910):

Solve for y: \[\Large x - 4y = 4\] \[\Large -4y = -x + 4\] \[\Large y = \frac{-x + 4}{-4}\] \[\Large y = \frac{-x}{-4} + \frac{4}{-4}\] \[\Large y = \frac{x}{4} -1\] \[\Large y = \frac{1}{4}x -1\] So the equation is now in y = mx+b form where m is the slope and b is the y-intercept Since m = 1/4, this means that the slope is 1/4. Since b = -1, the y-intercept is -1, which is the point (0,-1)